OpenStudy (anonymous):

find the inverse of f(x) = 1-x / x

OpenStudy (bahrom7893):

You can find the inverse like this: \[y=\frac{1-x}{x}\]\[xy=1-x\]\[xy+x=1\]\[x(y+1)=1\]\[x=\frac{1}{1+y}\]

OpenStudy (bahrom7893):

Now just replace y with x to get: \[f^{-1}(x)=\frac{1}{1+x}\]

OpenStudy (anonymous):

pls show the verification f(f^-1(x)) @bahrom7893

OpenStudy (bahrom7893):

To verify we need to confirm that \[f(f^{-1}(x))=1\]

OpenStudy (bahrom7893):

\[f(\frac{1}{1+x})=\frac{1-\frac{1}{1+x}}{\frac{1}{1+x}}\]

OpenStudy (bahrom7893):

Do the algebra and confirm that it eventually equals 1.

OpenStudy (anonymous):

ty